# Lifelong Learning with Behavior Consolidation for Vehicle Routing

### v2

@@ D.2 Main Experiments on TSP
Order Method k=3 k=6 AP AF AMF APl AG AP AF AMF APl AG 1 Restart 67.0 11.3 11.3 10.0 24.6 39.9 40.8 68 7.1 57.4 Fine-tuning 31.2 5.3 5.3 4.2 25.9 18.2 36.5 47.9 2.5 35.6 EWC 26.2 4.5 4.5 5.1 27.4 17.3 22.4 22.6 4.0 29.2 LiBOG 30.2 3.7 3.7 5.2 28.3 19.4 15.2 22.1 4.3 38.2 LLR-BC 2.9 0.3 0.3 2.5 19.1 1.7 0.8 0.9 1.3 21.6 2 Restart 39.3 10.0 10.0 5.4 68.7 29.7 128.4 128.4 7.1 76.0 Fine-tuning 16.6 7.9 7.9 2.4 61.3 13.8 42.5 42.5 2.7 59.3 EWC 15.4 8.5 8.5 4.2 66.1 14.5 18.9 19.2 4.7 48.1 LiBOG 18.3 8.0 8.0 4.0 63.0 18.7 16.8 16.8 4.7 45.1 LLR-BC 4.5 0.1 0.1 2.8 50.6 2.5 0.5 0.6 1.6 27.6 3 Restart 49.9 56.8 56.8 3.1 70.5 39.9 14.6 42.8 7.1 72.9 Fine-tuning 19.9 29.2 29.2 1.8 17.6 15.4 15.4 30.5 2.2 41.5 EWC 19.8 19.9 19.9 2.6 16.7 17.0 10.8 19.0 3.8 37.4 LiBOG 20.6 15.6 15.6 2.1 16.7 17.4 11.8 20.2 4.3 37.2 LLR-BC 2.2 0.0 0.0 1.9 10.0 3.0 0.6 0.9 1.8 22.9 4 Restart 40.9 46.0 46.0 7.0 156.2 29.7 37.6 47.5 7.1 85.7 Fine-tuning 25.1 32.2 32.2 7.4 164.3 15.5 33.2 33.2 4.4 85.2 EWC 34.1 34.1 34.1 8.6 160.8 22.8 23.2 31.2 7.0 77.6 LiBOG 32.4 34.4 34.4 8.7 154.5 22.2 25.5 33.1 7.5 74.5 LLR-BC 8.9 1.1 1.1 7.7 154.6 5.0 1.4 1.4 4.2 68.9 5 Restart 14.0 11.1 11.1 8.8 99.5 19.5 30.9 41.5 7.1 69.9 Fine-tuning 11.5 9.1 9.1 9.4 99.1 11.3 16.7 28.9 5.7 65.4 EWC 21.5 12.0 12.0 8.9 102.5 20.0 17.7 31.0 8.1 68.3 LiBOG 18.7 9.4 9.4 8.6 92.1 18.2 16.9 22.0 8.2 63.5 LLR-BC 7.5 0.8 0.8 9.1 106.3 4.8 0.5 1.7 5.3 64.5
@@
D.3 Result On Benchmark Instances
We use the identical representative instance set as selected and used by Zhou et al. (2023) Table 7 and 8 present the detailed test performance on CVRPLIB (Set-X) and TSPLIB instances, respectively. Following the same protocol of the main experiments, the reported values are normalized based on the base solution found in our experiments instead of the best-known/optimal solutions reported in existing studies.
@@ D.3 Result On Benchmark Instances
Instance Restart Fine-tuning EWC LiBOG LLR-BC X-n101-k25 8.16 17.92 0 10.56 8.93 X-n153-k22 38.03 0 21.71 43.18 23.72 X-n200-k36 19.54 8.13 7.02 13.22 0 X-n251-k28 14.26 0 14.25 9.23 1.47 X-n303-k21 66.62 10.03 38.12 27.47 0 X-n351-k40 118.54 5.53 6.5 18.25 0 X-n401-k29 23.33 1.48 16.77 21.85 0 X-n459-k26 76.71 0 9.8 8.78 20.85 X-n502-k39 9.86 0 4.42 21.88 20.08 X-n548-k50 35.99 0 9.31 6.18 4.33 X-n599-k92 27.25 17.92 7.76 25.96 0 X-n655-k131 31.02 26.97 0 11.66 0.77 X-n701-k44 58.96 0 30.82 9.88 13.01 X-n749-k98 43.41 20.44 0 13.76 7.94 X-n801-k40 29.82 0 2.3 1.55 14.34 X-n856-k95 50.49 29.5 13.87 35.03 0 X-n895-k37 71.45 0 33.18 26.84 21 X-n957-k87 58.04 24.37 9.06 27.86 0 X-n1001-k43 48.51 0 0.5 9.78 13.28 Mean (Std.) 43.68 (27.37) 8.54 (10.71) 11.86 (11.59) 18.05 (10.86) 7.88 (8.75)
@@
Instance Restart Fine-tuning EWC LiBOG LLR-BC kroA100 86.88 40.56 42.67 75.6 0 kroA150 29.43 28.06 0 7.58 12.15 kroA200 101.65 0 16.93 44.99 48.77 kroB200 105.78 10.86 34.53 12.47 0 ts225 152.45 60.52 70.33 24.24 0 tsp225 97.43 9.61 0 15 2.47 pr226 17.84 0 5.95 16.59 16.12 pr264 47.94 36.59 0 34.88 46.66 a280 72.55 8.98 0 28.86 19.19 pr299 79.27 0 5.09 6.53 16.29 lin318 97.15 0 16 46.94 32.34 rd400 81.68 18.25 0 25.14 20.14 fl417 40.69 3.67 30.88 0 21.12 pr439 126.81 29.1 3.8 0 0.14 pcb442 108.27 0 1.17 17.31 11.12 d493 253.56 443.1 280.94 123.82 0 u574 92.14 19.11 3.25 10.8 0 rat575 122.04 0 24.43 20.29 52.95 p654 45.81 0 11.12 5.04 26.84 d657 267.62 136.78 81.08 101.14 0 u724 66.06 30.1 0 2.83 25.28 rat783 135.86 0 10.48 23.03 34.46 pr1002 65.53 2.33 0.67 0 29.88 Mean (Std. ) 99.76(61.06) 38.16(93.36) 27.80(59.58) 27.96(32.29) 18.08(16.98)
D.4 Hyperparameter Sensitivity Analysis Details
As Table 3 demonstrated, on CVRP, increasing $\alpha$ reduces AF and AMF while increasing APl as expected, as it leads to greater emphasis on stability over plasticity, which aligns with expectations since a higher weight in the consolidation loss term will lead to more focus on stability rather than plasticity. In contrast, on the TSP, increasing $\alpha$ results in worse performance across evaluation metrics. A potential reason is that TSP tasks are more similar to each other than those in CVRP, as TSP does not involve differences in demand distribution. This higher similarity increases the likelihood of beneficial backward transfer when learning a new task. In such cases, using a lower consolidation weight could be more advantageous. Nevertheless, LLR-BC consistently outperforms all baseline methods under each tested setting (c.f., Table 1). No significant pattern in performance changing is found by varying $|\mathcal{E}|$ or $|\mathcal{B}|$. But the impact of changing them is substantially small, verifying the robustness of LLR-BC.
D.5 Details of Applicability Experiments

### v3

@@ D.2 Main Experiments on TSP
The average solution distance and optimality gap (optimal solutions are obtained by Gurobi) of the solver trained with each method on all tasks follwoing task order 1 are listed in Table 7. Notably, we use the solver trained after the corresponding task to evaluate Restart, making it single task method.
@@
Order Method k=3 k=6 AP AF AMF APl AG AP AF AMF APl AG 1 Li (inter) 17.5 0.0 0.0 14.8 22.5 20.2 1.4 1.4 18.9 22.7 Li (intra) 42.9 1.4 1.4 51.4 59.1 45.9 0.1 0.7 50.3 54.5 Feng 25.3 4.8 4.8 15.9 23.5 26.5 0.7 5.5 21.6 25.1 Restart 67.0 11.3 11.3 10.0 24.6 39.9 40.8 68 7.1 57.4 Fine-tuning 31.2 5.3 5.3 4.2 25.9 18.2 36.5 47.9 2.5 35.6 EWC 26.2 4.5 4.5 5.1 27.4 17.3 22.4 22.6 4.0 29.2 LiBOG 30.2 3.7 3.7 5.2 28.3 19.4 15.2 22.1 4.3 38.2 LLR-BC 2.9 0.3 0.3 2.5 19.1 1.7 0.8 0.9 1.3 21.6 2 Li (inter) 37.6 0.0 0.0 35.8 53.7 38.0 0.5 0.8 35.8 42.6 Li (intra) 58.1 0.6 0.6 59.7 92.8 55.3 0.7 1.0 56.6 69.7 Feng 29.5 1.1 1.1 25.1 56.4 26.6 2.0 2.0 22.7 35.5 Restart 39.3 10.0 10.0 5.4 68.7 29.7 128.4 128.4 7.1 76.0 Fine-tuning 16.6 7.9 7.9 2.4 61.3 13.8 42.5 42.5 2.7 59.3 EWC 15.4 8.5 8.5 4.2 66.1 14.5 18.9 19.2 4.7 48.1 LiBOG 18.3 8.0 8.0 4.0 63.0 18.7 16.8 16.8 4.7 45.1 LLR-BC 4.5 0.1 0.1 2.8 50.6 2.5 0.5 0.6 1.6 27.6 3 Li (inter) 6.1 0.0 0.0 6.9 9.0 19.6 0.3 0.3 19.6 22.8 Li (intra) 22.0 0.0 0.0 29.5 27.5 46.3 0.0 0.0 51.4 57.2 Feng 9.4 3.7 3.7 1.3 11.3 25.6 3.9 4.3 21.5 23.8 Restart 49.9 56.8 56.8 3.1 70.5 39.9 14.6 42.8 7.1 72.9 Fine-tuning 19.9 29.2 29.2 1.8 17.6 15.4 15.4 30.5 2.2 41.5 EWC 19.8 19.9 19.9 2.6 16.7 17.0 10.8 19.0 3.8 37.4 LiBOG 20.6 15.6 15.6 2.1 16.7 17.4 11.8 20.2 4.3 37.2 LLR-BC 2.2 0.0 0.0 1.9 10.0 3.0 0.6 0.9 1.8 22.9 4 Li (inter) 98.5 0.0 0.0 104.5 158.1 109.6 0.0 0.0 119.1 146.4 Li (intra) 73.9 0.4 0.4 100.2 203.6 63.7 0.0 0.1 79.1 122.6 Feng 20.6 0.0 0.0 21.1 186.7 17.9 0.0 1.2 18.5 84.6 Restart 40.9 46.0 46.0 7.0 156.2 29.7 37.6 47.5 7.1 85.7 Fine-tuning 25.1 32.2 32.2 7.4 164.3 15.5 33.2 33.2 4.4 85.2 EWC 34.1 34.1 34.1 8.6 160.8 22.8 23.2 31.2 7.0 77.6 LiBOG 32.4 34.4 34.4 8.7 154.5 22.2 25.5 33.1 7.5 74.5 LLR-BC 8.9 1.1 1.1 7.7 154.6 5.0 1.4 1.4 4.2 68.9 5 Li (inter) 75.1 0.1 0.1 100.5 161.7 95.2 0.0 0.1 115.0 153.6 Li (intra) 42.6 2.0 2.0 69.8 127.3 60.0 0.2 0.8 76.6 107.9 Feng 17.3 1.5 1.5 13.1 100.0 23.9 2.6 5.8 22.8 58.5 Restart 14.0 11.1 11.1 8.8 99.5 19.5 30.9 41.5 7.1 69.9 Fine-tuning 11.5 9.1 9.1 9.4 99.1 11.3 16.7 28.9 5.7 65.4 EWC 21.5 12.0 12.0 8.9 102.5 20.0 17.7 31.0 8.1 68.3 LiBOG 18.7 9.4 9.4 8.6 92.1 18.2 16.9 22.0 8.2 63.5 LLR-BC 7.5 0.8 0.8 9.1 106.3 4.8 0.5 1.7 5.3 64.5
@@
D.3 Comparison with Non-lifelong Settings
Notably, in lifelong learning scenarios, training tasks arise sequentially, and only the current task can be learned. Therefore, multi-task solvers, which learn from training tasks simultaneously and encounter no forgetting issue, are not applicable. To further enhance our experimental study, we implement 2 straightforward multi-task learning methods (batch-level (BL) and epoch-level (EL) task switching) based on POMO, denoted as POMO-MT-BL and POMO-MT-EL, and report the model’s performance after learn from all 6 tasks. Results, as follows, demonstrate that the difference is very small, and LLR-BC even outperforms the multi-task settings in the last 4 tasks of CVRP and all tasks of TSP. In addition, we add a comparison with INViT (the recent ancd high-performance, arguable state-of-the-art, cross-distribution and cross-scale neural solver) in the original setting without adaptation to lifelong learning. Specifically, we directly use a model provided by Fang et al. (2024), which is trained with a large budget on the uniform distribution scale 100 task and demonstrates good zero-shot generalization ability to unseen tasks. Comparison results indicate that LLR-BC, with lifelong learning, outperforms INViT that without requiring additional training or lifelong learning mechanisms. It further verifies the importance of lifelong learning and the effectiveness of our proposed LLR-BC. Detailed results are listed in Table 6 and 7.
D.4 Result On Benchmark Instances
We use the identical representative instance set as selected and used by Zhou et al. (2023) Table 9 and 10 present the detailed test performance on CVRPLIB (Set-X) and TSPLIB instances, respectively. Following the same protocol of the main experiments, the reported values are normalized based on the base solution found in our experiments instead of the best-known/optimal solutions reported in existing studies.
@@ D.4 Result On Benchmark Instances
Instance Li (intra) Li (inter) Feng Restart Fine-tuning EWC LiBOG LLR-BC X-n101-k25 38.33 19.09 42.64 8.16 17.92 0 10.56 8.93 X-n153-k22 52.41 12.65 30.32 38.03 0 21.71 43.18 23.72 X-n200-k36 26.28 21.48 10.85 19.54 8.13 7.02 13.22 0 X-n251-k28 12.71 12.98 14.37 14.26 0 14.25 9.23 1.47 X-n303-k21 83.59 44.96 18.77 66.62 10.03 38.12 27.47 0 X-n351-k40 43.88 65.26 0 118.54 5.53 6.5 18.25 0 X-n401-k29 29.12 39.31 30.61 23.33 1.48 16.77 21.85 0 X-n459-k26 81.01 60.34 66.94 76.71 0 9.8 8.78 20.85 X-n502-k39 32.57 87.95 31.46 9.86 0 4.42 21.88 20.08 X-n548-k50 25.36 6.86 5.78 35.99 0 9.31 6.18 4.33 X-n599-k92 55.1 25.09 6.62 27.25 17.92 7.76 25.96 0 X-n655-k131 66.91 60.46 15.64 31.02 26.97 0 11.66 0.77 X-n701-k44 34.08 14.51 21.06 58.96 0 30.82 9.88 13.01 X-n749-k98 57.64 46.36 41.03 43.41 20.44 0 13.76 7.94 X-n801-k40 17.03 24.97 20.79 29.82 0 2.3 1.55 14.34 X-n856-k95 21.53 34.56 47.96 50.49 29.5 13.87 35.03 0 X-n895-k37 119.88 53.02 17.84 71.45 0 33.18 26.84 21 X-n957-k87 57.95 74.26 61.91 58.04 24.37 9.06 27.86 0 X-n1001-k43 34.88 43.83 35.69 48.51 0 0.5 9.78 13.28 Mean (Std.) 46.86(26.95) 39.37(23.18) 27.38(18.53) 43.68 (27.37) 8.54 (10.71) 11.86 (11.59) 18.05 (10.86) 7.88 (8.75)
@@
Instance Li (intra) Li (inter) Feng Restart Fine-tuning EWC LiBOG LLR-BC kroA100 115.35 56.74 44.05 86.88 40.56 42.67 75.6 0 kroA150 85.3 56.6 11.67 29.43 28.06 0 7.58 12.15 kroA200 116.89 73.23 96.39 101.65 0 16.93 44.99 48.77 kroB200 90.12 46.27 55.42 105.78 10.86 34.53 12.47 0 ts225 145.3 27.3 102.23 152.45 60.52 70.33 24.24 0 tsp225 76.08 31.3 28.75 97.43 9.61 0 15 2.47 pr226 38.69 30.16 42.36 17.84 0 5.95 16.59 16.12 pr264 279.81 164.86 164.17 47.94 36.59 0 34.88 46.66 a280 119.41 43.49 80.83 72.55 8.98 0 28.86 19.19 pr299 127.96 64.42 65.24 79.27 0 5.09 6.53 16.29 lin318 127.91 61.94 104.45 97.15 0 16 46.94 32.34 rd400 128.69 53.32 55.73 81.68 18.25 0 25.14 20.14 fl417 28.2 58.18 80.35 40.69 3.67 30.88 0 21.12 pr439 120.37 50.77 61.96 126.81 29.1 3.8 0 0.14 pcb442 157.56 79.02 81.29 108.27 0 1.17 17.31 11.12 d493 40 52.33 89.3 253.56 443.1 280.94 123.82 0 u574 127.96 43.75 73.05 92.14 19.11 3.25 10.8 0 rat575 152.37 51.56 77.56 122.04 0 24.43 20.29 52.95 p654 150.62 56.15 60.28 45.81 0 11.12 5.04 26.84 d657 126.42 50.54 104.57 267.62 136.78 81.08 101.14 0 u724 161.09 58.65 83.95 66.06 30.1 0 2.83 25.28 rat783 155.97 65.08 75.42 135.86 0 10.48 23.03 34.46 pr1002 147.73 68.96 89.49 65.53 2.33 0.67 0 29.88 Mean (Std. ) 122.60(51.78) 58.46(26.55) 75.15(30.83) 99.76(61.06) 38.16(93.36) 27.80(59.58) 27.96(32.29) 18.08(16.98)
D.5 Hyperparameter Sensitivity Analysis Details
As Table 3 demonstrated, on CVRP, increasing $\alpha$ reduces AF and AMF while increasing APl as expected, as it leads to greater emphasis on stability over plasticity, which aligns with expectations since a higher weight in the consolidation loss term will lead to more focus on stability rather than plasticity. In contrast, on the TSP, increasing $\alpha$ results in worse performance across evaluation metrics. A potential reason is that TSP tasks are more similar to each other than those in CVRP, as TSP does not involve differences in demand distribution. This higher similarity increases the likelihood of beneficial backward transfer when learning a new task. In such cases, using a lower consolidation weight could be more advantageous. Nevertheless, LLR-BC consistently outperforms all baseline methods under each tested setting (cf. Table 1). No significant pattern in performance changing is found by varying $|\mathcal{E}|$ or $|\mathcal{B}|$. But the impact of changing them is substantially small, verifying the robustness of LLR-BC.
D.6 Details of Applicability Experiments
@@
Appendix E LLM Usage Statement
We used ChatGPT (GPT-5) only as an assistive tool for grammar checking and language polishing. The model was not involved in research ideation, algorithm design, experiment execution, or result analysis. All scientific content and conclusions are entirely the work of the authors.